OpenStudy (anonymous):

u2 - (2)(8) + 82 u squared minus 2 times 8 + 82. Can someone help figure out how to factor this further?

OpenStudy (anonymous):

Are you sure your question is right ???

OpenStudy (anonymous):

I think you are missing \(u\) somewhere in your question...

OpenStudy (anonymous):

Is not it like this : \[u^2 - (2)(8)u +82\]

OpenStudy (anonymous):

u2 - 16u + 64 = u2 - (2)(8) + 82 they are saying the answer is = (u - 8)2 or (u - 8) squared U squared minus 16U + 64 = U squared - 2 times 8 + 8 squared

OpenStudy (anonymous):

I am not understanding this at all

OpenStudy (anonymous):

Sorry I cant figure out how to add equations thanks anyway

OpenStudy (anonymous):

\[u^{2} - 16u + 64 \implies u^2 - 8u - 8u + 64 \implies u(u-8) - 8(u - 8)\] \[\implies \color{blue}{(u-8)^2}\]

OpenStudy (anonymous):

Gotta go....